OpenStudy (jennychan12):

If F(x) = integral from x to 1 f(t)dt, where f(t) = integral from x^2 to 1 of [(sqrt(10+u^4))/u]du, find F''(2).

OpenStudy (jennychan12):

If F(x) = \[\int\limits_{1}^{x}f(t)dt\] where f(t) = \[\int\limits_{1}^{t^2}(\frac{ \sqrt{10+u^4} }{ u })du\] find F''(2)